# Observing the Sun with the Atacama Large Aperture Submillimeter Telescope (AtLAST): Forecasting Full-disk Observations

### v1

@@ 4.2.3 Scanning strategies for large FOV instruments (Case C)
The double-circle pattern has proved to be an efficient method of scanning the solar disk when utilising instruments with small to intermediately sized fields of view. However, as one gets to larger FOV instruments such a method would prove less efficient. With very large FOV instruments, one can get away with simpler scan pattern. Assuming the instrument FOV diameter to be greater than 1 R⊙, simply scanning in a circular motion would be sufficient to observe the full solar disk. Such an instrument would then be utilised in circular scans centred on the solar disk with a radius of half the instrument FOV, and thusly cover the full disk on possibly very short time scales. Such a case is considered, and we here consider the case where one wishes for the entire solar disk and off-limb to be fully sampled, meaning that the area we want to scan has a diameter of 2.4 R. Using an instrument with a FOV of 1.2 R⊙ and scanning along a circle of radius 0.6 R⊙ will then cover the full disk, assuming the sample rate is sufficiently high and that the instrument FOV is sufficiently filled with detector elements to not leave gaps. An advantage of this scanning method is its simplicity, as it requires very little stress on the telescope, as the acceleration is kept low and slowly varying.
@@ 4.3 Cadence of observations
To arrive at such a relation, a double-circle scanning pattern was initially assumed, as the number of minor circles per pattern could directly be related to both the needed scan time to complete such a pattern, but also instrument FOV. Estimating the number of minor circles required to fully sample a region of a given size with a specified instrument could therefore yield a relation between the total needed scan time and the instrument FOV. To cover each point in a sky patch at least once, it is sufficient that the minor circles in the double-circle pattern are spaced out in such a way that gaps in-between them correspond to the diameter of the instrument FOV. The full region in-between two minor circles would thusly be sampled fully by two consecutive minor circle paths. The gap in-between two consecutive minor circles is called the sampling length (White et al., 2017), and in the case described here is equal to the diameter of the instrument’s FOV, FOV. Lowering the sampling length to FOV fully samples all regions in-between two consecutive minor circles twice.
To find a relation between the instrument’s instantaneous FOV and the total scan time, the size of the sky patch was set to be a circular region of diameter 2400 ”, the same size as the region scanned by ALMA TP antennas when mapping the solar disk (White et al., 2017). It was chosen to cover the full disk and a region of over 200 ” off-limb, advantageous both for calibration and for observing solar prominences (Gunár et al., 2016, 2018; White et al., 2017). A wide range of instrument FOVs were then set-up, and the required number of minor circles per pattern computed, both with sampling lengths of one and a half the diameter of the instrument FOV. This resulted in the relations:
@@
As discussed in Sect. 4.1.1, the instrument’s instantaneous FOV is itself dependent on several instrumental properties, here the observing frequency, the number of pixels and the pixel spacing. It is therefore possible to instead relate the scan time to the number of pixels at set observing frequencies and with specified pixel spacings. However, note that the assumption of the full instantaneous FOV being sufficiently filled with detectors to instantaneously cover the sky region it passes over still applies. This only applies to setups where we have sufficient numbers of detector elements to not leave gaps after the full detector array has passed the sky patch. We therefore do not expect any found trends to be applicable at large detector spacings when the number of detectors is low. However, we argue that the relations are still valid at detector counts currently considered for AtLAST. As a first generation multi-chroic camera at AtLAST is expected to have four frequency channels, we do this for four different frequencies, and for the three pixel spacings 0.5, 1.0, and 2.0 f for a sampling length of FOV. The four frequency channels were chosen to be centred on 100, 250, 670, and 950 GHz. The three first bands are then counterparts to Band 3, 6, and 9 at ALMA222https://www.eso.org/public/teles-instr/alma/receiver-bands/ , while the last is meant to check the feasibility of using AtLAST for high-cadence and high-frequency observations of the full disk.
The relations between the scan time and the number of pixels are shown in Fig. 6, where black dashed lines show the expected maximum number of pixels for a 1st and 2nd generation multi-chroic instrument (van Kampen, 2024). Blue, red, and green are here used to show different pixel spacings. One can here see the large effect multi-pixel instruments can have on the cadence of full-disk solar observations, as even going to the 50,000 pixel first generation instrument results in sub-minute cadence across all the four frequency bands, except in 670 and 950 GHz with a pixel spacing of 0.5 f. The improvement in cadence over ALMA TP full-disk maps is then on the order of ten, in addition to a four times higher angular resolution.
@@ 5.1 Required cadence
The trends in scan time as a function of detector counts for the four frequency bands considered in Sect. 4.3, and shown in Fig. 6, were made with the scan pattern spaced out in such a way as to ensure each patch of sky is passed at least twice by the instrument, i.e. a sampling length of half the FOV diameter of the utilised instrument ( FOV). Without more detailed knowledge of the time scales needed to observe the shortest lived dynamic phenomena in the chromosphere, it is difficult to decide on a lower limit cadence and thus, via the relation between cadence and pixel count, the minimum number of pixels required for a set pixel spacing at each observing frequency. It should be noted that typical periods of quasi-periodic pulsations that occur during flares range from a fraction of a second to sev- eral minutes (see, e.g., Parks and Winckler, 1969; Wedemeyer et al., 2016, and references therein).
The number of pixels required for the considered frequency bands at different pixel spacings to reach a one minute cadence are given in Table. 2. We find that the number of required pixels is highly dependent on both the observing frequency and the spacing between pixels, as expected. A first generation multi-chroic camera with 50,000 pixels per frequency band, in accordance to the specifications as provided by the AtLAST instrumentation group (van Kampen, 2024), could achieve a cadence of 1 min across the considered frequency bands and pixel spacings (see Table 2), except with a spacing of f observing at 950 GHz. Interestingly, less than 25,000 pixels are required to observe up to 950 GHz with a spacing of 1 f, well within the estimates of a 1st generation instrument while still achieving a high cadence. This is encouraging for an instrument that employs dedicated (monochroic) detectors for different frequency ranges. The detector for the highest covered frequency range requires the highest number of pixels, thus setting the limit for the overall instrument. On the other hand, the detectors for the lower frequencies would require fewer pixels while still able to reach a 1-min-cadence.
@@ 5.2 Efficient scan patterns
We have seen that both the Lissajous daisy and double-circle scan patterns are suitable for efficiently scanning sky patches, at least with instruments of reasonably small FOVs. As seen for Case C (see Sect. 4.2.3), when the FOV of the instrument becomes sufficiently large, scanning in an extended and often complex scan pattern is no longer as efficient. Instead, a simple scan pattern like a simple circular path is sufficient. The point where the FOV becomes sufficiently large as to warrant a change of scan strategy depends only on the size of the sky region of interest. As for the circular region of radius ” considered in Sect. 4.3, corresponding to the extent of the solar disk and a 200” off-limb region as observed by ALMA (Phillips et al., 2015), a circular instrument with an FOV diameter of 1200” would be large enough to justify the use of a simple circular scan. The scan time for this FOV is independent of frequency, as long as the FOV is sufficiently filled with detectors at the respective frequencies.
Assuming monochroic detectors, this results in much fewer pixels being required to fill the sufficiently large FOV to make the double-circle pattern inefficient at lower frequencies as compared to higher frequencies. Table 3 shows the number of required detectors to fill the 1200” diameter FOV which could be used in single-circle scans at different observing frequencies and detector spacings. It is here found that, if we want to push for observing at 950 GHz and utilise a single-circle scan strategy, detectors spacings of less than two would not be viable, as they would require detector counts above the current estimate for even a second generation instrument at AtLAST (¡ 300,000, van Kampen, 2024).
Observing Frequency | |||||
100 GHz | 250 GHz | 670 GHz | 950 GHz | ||
0.5 f | 21,080 | 131,645 | 945,622 | 1,901,240 | |
1.0 f | 5,279 | 32,943 | 236,420 | 475,305 | |
2.0 f | 1,320 | 8,223 | 59,108 | 118,853 | |
Detector spacing |
@@ 5.3 Scan coverage
A Lissajous daisy, as used by the Green Bank Telescope (GBT) (Mason, 2003; Korngut et al., 2011), has the advantage of a near-uniformly sampled central region when applying a centre-offset. However, the uniformly sampled region is often small compared to the size of the full scan pattern, meaning that the total scan path is long, requiring long scan times to achieve a uniformly sampled central region large enough to cover the solar disk. Nevertheless, the Lissajous daisy is an option that should be explored, in combination with other scan strategies. However, the suitability of the considered scan strategiesfor solar observing with AtLAST will depend on the details of the employed instrument. When considering large instruments, it is expected that the scan pattern has less of an impact on both the image sampling and the resulting cadence, as one can fill the solar disk on short time scales simply by following a simple and often short scan path. However, for smaller instruments, as seen in Sect. 4.2.1, a larger and more complex scan strategy is required to fully sample the solar disk, thus making the choice of a suitably efficient scan path all the more important. This both affects the sampling of the source, but also the resulting cadence.
We here compare four scan strategies that could prove relevant for solar observing at AtLAST, namely the double circle, the Lissajous daisy, the raster, and the Lissajous box. For the comparison of these scan strategies, we consider a small instrument with 1000 pixels observing at 950 GHz. With a detector spacing of 2 f this corresponds to a FOV of . The resulting tight scan patterns are shown in the left-most row in Fig. 8. Note that the scan time for fully sampling the solar disk varies greatly between the strategies: 95 s for the double-circle, 102 s for the back-and-forth raster, 142 s for the Lissajous box, and 250 s for the Lissajous daisy pattern. The reason for the relatively long scan time needed for the Lissajous daisy to fully sample the whole solar disk is its tendency to miss small patches on the disk, so that it takes long before all patches have been covered. We conclude that the other three scan patterns provide a more even and thus efficient sampling. This is especially true fpr the back-and-forth raster and the double-circle scanning the disk either from top to bottom or radially. These two patterns also have the advantage that in a possible malfunction that halts the scanning, the mid-scan data would still be usable, as they cover connected and fully sampled regions on the Sun.
@@ 5.4 Quantifying the quality of synthetic observations
In the top-row of Figure 11, the spatial power spectra are shown for each of the three cases A, B, and C, for the available observing frequencies (i.e. at 100, 250, 670, and 950 GHz for A and C, and for 100, 250, and 670 GHz for Case B). See Figure 3 for the corresponding input maps. The power spectra show greatly reduced power at smaller scales for all synthetic observations compared to the input map, which is observed across all frequencies and with all three cases. The maps that required post-processing in the form of smoothing, i.e., the maps at 100 GHz for cases A and C, and the 250 GHz map for case C, show no signe of a plateau at or near the spatial scale of beam at those frequencies. All other maps show a flattening of the power spectra at spatial scales just larger than the beam scale.
@@ 6 Conclusions
Based on our findings that high cadence full-disk solar observations with AtLAST are possible, we recommend that a dedicated instrumental study be performed for finding the optimal parameters of a solar- or at least solar-capable instrument for AtLAST, which would undoubtedly provide the solar millimetre community with extremely useful data.
@@ Acknowledgements
This work was carried out in connection with the AtLAST design study, which has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No. 951815, and from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101188037 (AtLAST2). This work is supported by the Research Council of Norway through its Centres of Excellence scheme, project number 262622 (Rosseland Centre for Solar Physics). This paper makes use of the following ALMA data: ADS/JAO.ALMA#2022.1.01544.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC(Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in co-operation with the Republic of Chile. The authors thank Evanthia Hatziminaoglou and Matt Griffin for their valuable comments.

### v2

@@ 4.2.3 Scanning strategies for large FOV instruments (Case C)
The double-circle pattern has proved to be an efficient method of scanning the solar disk when utilising instruments with small to intermediately sized fields of view. However, as one gets to larger FOV instruments such a method would prove less efficient. With very large FOV instruments, one can get away with simpler scan pattern.
Assuming the instrument FOV diameter to be greater than 1 R⊙, simply scanning in a circular motion would be sufficient to observe the full solar disk. Such an instrument would then be utilised in circular scans centred on the solar disk with a radius of half the instrument FOV, and thusly cover the full disk on possibly very short time scales. Such a case is considered, and we here consider the case where one wishes for the entire solar disk and off-limb to be fully sampled, meaning that the area we want to scan has a diameter of 2.4 R. Using an instrument with a FOV of 1.2 R⊙ and scanning along a circle of radius 0.6 R⊙ will then cover the full disk, assuming the sample rate is sufficiently high and that the instrument FOV is sufficiently filled with detector elements to not leave gaps. An advantage of this scanning method is its simplicity, as it requires very little stress on the telescope, as the acceleration is kept low and slowly varying.
@@ 4.3 Cadence of observations
To arrive at such a relation, a double-circle scanning pattern was initially assumed, as the number of minor circles per pattern could directly be related to both the needed scan time to complete such a pattern, but also instrument FOV. Estimating the number of minor circles required to fully sample a region of a given size with a specified instrument could therefore yield a relation between the total needed scan time and the instrument FOV. To cover each point in a sky patch at least once, it is sufficient that the minor circles in the double-circle pattern are spaced out in such a way that gaps in-between them correspond to the diameter of the instrument FOV. The full region in-between two minor circles would thusly be sampled fully by two consecutive minor circle paths. The gap in-between two consecutive minor circles is called the sampling length (White et al., 2017), and in the case described here is equal to the diameter of the instrument’s FOV, FOV. Lowering the sampling length to FOV means that regions in-between two consecutive minor circles are covered twice, ensuring better sampling. This could be especially useful if the detector array is sparsely filled with pixels (detector elements).
To find a relation between the instrument’s instantaneous FOV and the total scan time, the size of the sky patch was set to be a circular region of diameter 2400 , the same size as the region scanned by ALMA TP antennas when mapping the solar disk (White et al., 2017). It was chosen to cover the full disk and a region of over 200 off-limb, advantageous both for calibration and for observing solar prominences (Gunár et al., 2016, 2018; White et al., 2017). A wide range of instrument FOVs were then set-up, and the required number of minor circles per pattern computed, both with sampling lengths of one and a half the diameter of the instrument FOV. This resulted in the relations:
@@
As discussed in Sect. 4.1.1, the instrument’s instantaneous FOV is itself dependent on several instrumental properties, here the observing frequency, the number of pixels and the pixel spacing. It is therefore possible to instead relate the scan time to the number of pixels at set observing frequencies and with specified pixel spacings. However, note that the assumption of the full instantaneous FOV being sufficiently filled with detectors to instantaneously cover the sky region it passes over still applies. This only applies to setups where we have sufficient numbers of detector elements to not leave gaps after the full detector array has passed the sky patch. We therefore do not expect any found trends to be applicable at large detector spacings when the number of detectors is low. However, we argue that the relations are still valid at detector counts currently considered for AtLAST. As a first generation multi-chroic camera at AtLAST is expected to have four frequency channels, we do this for four different frequencies, and for the three pixel spacings 0.5, 1.0, and 2.0 f for sampling lengths of and FOV. The four frequency channels were chosen to be centred on 100, 250, 670, and 950 GHz. The three first bands are then counterparts to Band 3, 6, and 9 at ALMA 222https://www.eso.org/public/teles-instr/alma/receiver-bands/, while the last is meant to check the feasibility of using AtLAST for high-cadence and high-frequency observations of the full disk.
The relations between the scan time and the number of pixels are shown in Fig. 6, where black dashed lines show the expected maximum number of pixels for a 1st and 2nd generation multi-chroic instrument (van Kampen, 2024). Blue, red, and green are here used to show different pixel spacings. One can here see the large effect multi-pixel instruments can have on the cadence of full-disk solar observations, as even going to the 50,000 pixel first generation instrument results in sub-minute cadence across all the four frequency bands, except in 670 and 950 GHz with a pixel spacing of 0.5 f. The improvement in cadence over ALMA TP full-disk maps is then on the order of ten, in addition to a four times higher angular resolution. It is also found that scanning at twice the sampling length (i.e., instead of FOV) impacts the scan time the same way as keeping the sampling length the same and doubling the spacing in-between each pixel in the instrument (i.e., instead of f). Many of the data points are therefore superimposed on each other, making them difficult to distinguish. To make this easier, the area between first-order log-log fits to the two different sampling lengths has been shaded to the appropriate colour associated with the pixel spacing.
@@ 5.1 Required cadence
The trends in scan time as a function of detector counts for the four frequency bands considered in Sect. 4.3, and shown in Fig. 6, were made with the scan pattern spaced out in such a way as to ensure each patch of sky is passed at least twice by the instrument, i.e. a sampling length of half the FOV diameter of the utilised instrument ( FOV). Without more detailed knowledge of the time scales needed to observe the shortest lived dynamic phenomena in the chromosphere, it is difficult to decide on a lower limit cadence and thus, via the relation between cadence and pixel count, the minimum number of pixels required for a set pixel spacing at each observing frequency. It should be noted that typical periods of quasi-periodic pulsations that occur during flares range from a fraction of a second to several minutes (see, e.g., Parks and Winckler, 1969; Wedemeyer et al., 2016, and references therein).
The number of pixels required for the considered frequency bands at different pixel spacings to reach a one minute cadence are given in Table. 2. We find that the number of required pixels is highly dependent on both the observing frequency and the spacing between pixels, as expected. A first generation multi-chroic camera with 50,000 pixels per frequency band, in accordance to the specifications as provided by the AtLAST instrumentation group (van Kampen, 2024), could achieve a cadence of 1 min across the considered frequency bands and pixel spacings (see Table 2), except with a spacing of f observing at 950 GHz. Interestingly, less than 25,000 pixels are required to observe up to 950 GHz with a spacing of 1 f, well within the estimates of a 1st generation instrument while still achieving a high cadence.
If one instead considers an instrument consisting of separate detector arrays of monochroic pixels for each frequency band, we find that all detector arrays could be at a f spacing except for 950 GHz, where the pixels would have to be spaced further apart ( f) for the required number of pixels to be within the estimate for a 1st generation instrument.
@@ 5.2 Efficient scan patterns
We have seen that the double-circle scan pattern can be suitable for efficiently scanning sky patches, at least with instruments of reasonably small FOVs. As seen for Case C (see Sect. 4.2.3), when the FOV of the instrument becomes sufficiently large, scanning in an extended and often complex scan pattern is no longer as efficient. Instead, a simple scan pattern like a simple circular path is sufficient. The point where the FOV becomes sufficiently large as to warrant a change of scan strategy depends only on the size of the sky region of interest. As for the circular region of radius considered in Sect. 4.3, corresponding to the extent of the solar disk and a 200 off-limb region as observed by ALMA (Phillips et al., 2015), a circular instrument with an FOV diameter of 1200 would be large enough to justify the use of a simple circular scan. The scan time for this FOV is independent of frequency, as long as the FOV is sufficiently filled with detectors at the respective frequencies.
As seen in Table. 3, assuming seperate monochroic detector arrays for each frequency, this results in much fewer pixels being required to fill the sufficiently large FOV to make the double-circle pattern inefficient at lower frequencies as compared to higher frequencies. Table 3 shows the number of required detectors to fill the 1200 diameter FOV which could be used in single-circle scans at different observing frequencies and detector spacings. It is here found that, if we want to push for observing at 950 GHz and utilise a single-circle scan strategy, detector spacings of less than f would not be viable, as they would require detector counts above the current estimate for even a second generation instrument at AtLAST ( 300,000, van Kampen, 2024).
| Obs. Freq. | 100 GHz | 250 GHz | 670 GHz | 950 GHz |
| Pixel. Spacing | ||||
| 0.5 f | 21 080 | 131 645 | 945 622 | 1 901 240 |
| 1.0 f | 5 279 | 32 943 | 236 420 | 475 305 |
| 2.0 f | 1 320 | 8 223 | 59 108 | 118 853 |
@@ 5.3 Scan coverage
A Lissajous daisy, as used by the Green Bank Telescope (GBT) (Mason, 2003; Korngut et al., 2011), has the advantage of a near-uniformly sampled central region when applying a centre-offset. However, the uniformly sampled region is often small compared to the size of the full scan pattern, meaning that the total scan path is long, requiring long scan times to achieve a uniformly sampled central region large enough to cover the solar disk. Nevertheless, the Lissajous daisy is an option that should be explored, in combination with other scan strategies. However, the suitability of the considered scan strategies for solar observing with AtLAST will depend on the details of the employed instrument.
When considering instruments with large and filled FOVs, it is expected that the scan pattern has less of an impact on both the image sampling and the resulting cadence, as one can fill the solar disk on short time scales simply by following a simple and often short scan path. However, for smaller instruments, as seen in Sect. 4.2.1, a larger and more complex scan strategy is required to fully sample the solar disk, thus making the choice of a suitably efficient scan path all the more important. This both affects the sampling of the source, but also the resulting cadence.
We here compare four scan strategies that could prove relevant for solar observing at AtLAST, namely the double circle, the Lissajous daisy, the raster, and the Lissajous box. For the comparison of these scan strategies, we consider a small instrument with 1000 pixels observing at 950 GHz. With a detector spacing of 2 f this corresponds to a FOV of . The resulting tight scan patterns are shown in the left-most row in Fig. 8. Note that the scan time for fully sampling the solar disk varies greatly between the strategies: 89 s for the double-circle, 102 s for the back-and-forth raster, 142 s for the Lissajous box, and 250 s for the Lissajous daisy pattern. The reason for the relatively long scan time needed for the Lissajous daisy to fully sample the whole solar disk is its tendency to miss small patches on the disk, so that it takes long before all patches have been covered. We conclude that the other three scan patterns provide a more even and thus efficient sampling. This is especially true for the back-and-forth raster and the double-circle scanning the disk either from top to bottom or radially. These two patterns also have the advantage that in a possible malfunction that halts the scanning, the mid-scan data would still be usable, as they cover connected and fully sampled regions on the Sun.
@@ 5.4 Quantifying the quality of synthetic observations
In the top-row of Figure 11, the spatial power spectra are shown for each of the three cases A, B, and C, for the available observing frequencies (i.e. at 100, 250, 670, and 950 GHz for A and C, and for 100, 250, and 670 GHz for Case B). See Figure 3 for the corresponding input maps. The power spectra show greatly reduced power at smaller scales for all synthetic observations compared to the input map, which is observed across all frequencies and with all three cases. The maps that required post-processing in the form of smoothing, i.e., the maps at 100 GHz for cases A and C, and the 250 GHz map for case C, show no sign of a plateau at or near the spatial scale of beam at those frequencies. All other maps show a flattening of the power spectra at spatial scales just larger than the beam scale.
@@ 6 Conclusions
Based on our findings that high cadence full-disk solar observations with AtLAST are possible, we recommend that a dedicated instrumental study be performed for finding the optimal parameters of a solar- or at least solar-capable instrument for AtLAST, which would undoubtedly provide the solar millimetre community with extremely useful data. Such an instrument, although optimised for high-cadence full-disk scanning of the Sun could also be used for observing other extended sources or sky regions. In such cases, the scanning speed and time could be altered based on the required integration time to observe the source in question.
@@ Acknowledgements
This project has received funding from the European Union’s Horizon Europe and Horizon2020 research and innovation programmes under grant agreements No. 101188037 (AtLAST2) and No. 951815 (AtLAST). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Research Executive Agency. Neither the European Union nor the European Research Executive Agency can be held responsible for them. This work is supported by the Research Council of Norway through its Centres of Excellence scheme, project number 262622 (Rosseland Centre for Solar Physics). This paper makes use of the following ALMA data: ADS/JAO.ALMA#2022.1.01544.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC(Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in co-operation with the Republic of Chile. The authors thank Evanthia Hatziminaoglou and Matt Griffin for their valuable comments.